50% of some number is 7 minus the same number
Answer by ptaylor(2198)

Let x=the number
50% of x=0.50x
7 minus the number is 7-x
so:
0.50x=7-x add x to each side
1.50x=7
x= 4 2/3 or 14/3
CK
50%=1/2 ok?
(1/2)(14/3)=7-(14/3)
7/3 =(21/3)-(14/3)
7/3 = 7/3
Hope this helps---ptaylor